OpenStudy (anonymous):

If f(x) = (3+x)/(x-3) what is f(a+2) A. (3+a)/(a-3) +2 B. (5+a).(a-1) C. 3+f(a+2)/f(a)-1

OpenStudy (anonymous):

A or c

OpenStudy (anonymous):

What do you think

OpenStudy (johnweldon1993):

Notice how it goes from \(\large f(x)\) to \(\large f(a + 2)\) This means take your original function which has 'x' in it...and replace every 'x' you see with (a + 2) \[\large f(x) = \frac{3+x}{x-3}\] \[\large f(a+2) = ? \]